# An Infinite Transitivity Theorem

### v3

@@ 4. Existence of P-states
We begin by fixing $\varphi_{n}\in P(A)$ s.t. $\varphi_{n}(f_{n})=1$. Since $(e_{n})$ strictly increasing and consists of projections, $\|f_{n}\|=1$ so such a state always exists. By [MU90, 3.3.9], these each have unique extensions $\tilde{\varphi_{n}}\in P(\mathscr{M}(A))$. Now define $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow S(\mathscr{Q}(A))$ by
$\varphi_{\mathscr{U}}(\pi(a))=\lim_{n\rightarrow\mathscr{U}}\tilde{\varphi_{n}}(a).$
We know these are states because they lift to weak-limits of pure states on $\mathscr{M}(A)$ and they are pure by [AN79, 1].
By theorem 2.8, it suffices to show that $\mathscr{F}_{\varphi_{\mathscr{U}}}$ is a quantum P-point iff $\mathscr{U}$ is a P-point. Suppose that $\mathscr{U}$ is a P-point and fix a sequence $(a_{n})$ in $\mathscr{F}_{\varphi}$ with respective lifts $\tilde{a_{n}}$. For each $m,n$, define
$X_{m,n}=\Big\{k\in\mathbb{N}:\tilde{\varphi_{k}}(\tilde{a_{n}})>1-\frac{1}{m}\Big\}.$
By construction, $X_{m,n}\in\mathscr{U}$. Since $\mathscr{U}$ is a P-point, there exists $X\in\mathscr{U}$ s.t. for all $m,n$, $X\setminus X_{m,n}$ finite. Enumerate $X=\{k_{j}:j\in\mathbb{N}\}$ so that for all $m,n$, $\tilde{\varphi_{k_{j}}}(\tilde{a_{n}})\geq 1-\frac{1}{m}$ for eventually all $j$. Therefore $\lim_{j\rightarrow\infty}\tilde{\varphi_{k_{j}}}(\tilde{a_{n}})=1$. Our claim is that $a=\pi(p_{X})$ is a lower bound for $(a_{n})$. Since $X\in\mathscr{U}$, $\varphi_{\mathscr{U}}(\pi(p_{X}))=1$, so $a\in\mathscr{F}_{\varphi}$. Now notice that
$\|\pi((\tilde{a_{n}}-1)p_{Y})\|=\lim_{k\rightarrow\infty}\|(\tilde{a_{n}}-1)p_{X}(1-e_{k})\|=\|(\tilde{a_{n}}-1)(p_{X}-p_{X})\|=0,$
since $p_{X}e_{n}\rightarrow p_{X}$ strictly. Therefore
$a_{n}a-a=\pi(\tilde{a_{n}}p_{X}-p_{X})=\pi((\tilde{a_{n}}-1)p_{X})=0,$
i.e. $a\ll a_{n}$ for all $n$.
Now suppose $\mathscr{U}$ is not a P-point, and fix decreasing $X_{n}\in\mathscr{U}$ with no pseudointersection in $\mathscr{U}$. Our claim is that $\pi(p_{X_{n}})$ has no lower bound in $\mathscr{P}_{\varphi_{\mathscr{U}}}$. By 4.2, any lower bound $a\in\mathscr{F}_{\varphi}$ of $(\pi(p_{X_{n}}))$ must in fact have $a\ll\pi(p_{X})$, where $X$ is a pseudointersection of $(X_{n})$, therefore no such $a$ exists. ∎
@@
Let $A$ be a nonunital $C^{*}$-algebra with a $\sigma$-unit of projections, and $\varphi\in P(\mathscr{Q}(A))$. If $\varphi$ is a P-state, then $\mathscr{X}_{\varphi}$ is a non-meagre P-filter. Also, for any decreasing $X_{n}\subseteq\mathbb{N}$, there exists a pseudointersection $X$ of $(X_{n})$ s.t. $\varphi(\pi(p_{X}))=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Also, if $\varphi(\pi(p_{X_{n}}))\not\rightarrow 0$, there exists $(k_{n})$ in $\mathbb{N}$ increasing and $X\subseteq\mathbb{N}$ s.t. $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ and $\varphi(X)>0$.
@@
Using the obvious type, there exists $a\in\mathscr{P}_{\varphi}$ s.t. $a\ll\pi(p_{X_{n}})$ and $\varphi(a)=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Then by 4.2, there is a pseudointersection $X$ of $(X_{n})$ s.t. $a\ll\pi(p_{X})$, so $1=\varphi(a)\leq\varphi(\pi(p_{X}))$. It then follows that $\mathscr{X}_{\varphi}$ is a P-filter.
@@
for $i<l$, where $k<m_{0}<\cdots<m_{l}$, then $l\epsilon<3^{n+1}$. We may then choose $k_{n}$ s.t. $\varphi(\pi(p_{X_{j}}))>\epsilon$ for some $k_{n}\leq j<k_{n+1}$ and
@@
Then by normality, we have $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ so that $\varphi(\pi(p_{X}))\geq\frac{\epsilon}{2}>0$. ∎
We now have enough to produce P-states.
@@
    For every nonunital $A$ with a $\sigma$-unit of projections, there exist $2^{\mathfrak{c}}$-many inequivalent P-states $\varphi$ on $\mathscr{Q}(A)$.
@@
    For every nonunital $A$ with a $\sigma$-unit of projections, there exist no P-states on $\mathscr{Q}(A)$.
@@
For (1), under the continuum hypothesis, there are $2^{\mathfrak{c}}$ RK-inequivalent selective ultrafilters (which are P-points) on $\mathbb{N}$, so the map $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ in 4.3 produces $2^{\mathfrak{c}}$ inequivalent P-states.
For (2), we use the forcing argument outlined in lemma [BI13, 6.5], where we use 4.4 in place of [BI13, 6.2]. In this model, there are no $\varphi$ in $\mathscr{Q}(A)$ having the conclusion of 4.4, so there are no P-states. ∎
@@ 4.1. Counterexample to Infinite Transitivity
The question of whether infinite transitivity in $(\mathscr{Q}(H),\pi_{\varphi})$ implies degree-1 saturation of $(\mathscr{Q}(H),\varphi)$ is fairly opaque. In this section, we at least prove that in ZFC, many massive coronas have an irreducible representation which does not admit infinite transitivity.
@@
For $A\subseteq\mathbb{N}$, $j\in\mathbb{N}$, we define $X+j=\{m+j:m\in X\}$ and for an ultrafilter $\mathscr{U}$ on $\mathbb{N}$, $\mathscr{U}+j=\{X+j:X\in\mathscr{U}\}$. Since $\beta\mathbb{N}$ is compact, we can define $\mathscr{U}+\mathscr{V}=\lim_{j\rightarrow\mathscr{V}}(\mathscr{U}+j)$. We call a nonprincipal ultrafilter $\mathscr{U}$ idempotent if $\mathscr{U}+\mathscr{U}=\mathscr{U}$.
By [EL58, 1], every compact topological space with a semigroup structure whose left operation is continuous admits an idempotent element. The addition defined in 4.6 plainly grants this semigroup structure on $\beta\mathbb{N}$, so there is $\mathscr{U}\in\beta\mathbb{N}$ idempotent, and this cannot be principal.
@@
Since we already proved the existence of $P$-states on $\mathscr{Q}(H)$ under the continuum hypothesis, it suffices to show there is an irreducible representation of $\mathscr{Q}(H)$ which fails infinite transitivity. In fact, such a representation exists in ZFC.
Let $\mathscr{U}$ be an idempotent ultrafilter on $\mathbb{N}$. Let $s\in\mathscr{B}(H)$ be the right shift along an orthonormal basis $(\eta_{n})$ of $H$ and $\varphi_{\mathscr{U}}$ as in 4.3 along $(\eta_{n})$. Let $\xi_{j}=\pi(s^{j})+L_{\varphi_{\mathscr{U}}}$. Then the equations
$\pi_{\varphi}(a)\xi_{j}=\delta_{j0}\xi_{j}$
@@
where $a_{n}=\langle a\eta_{n},\eta_{n}\rangle$. Thus $X_{1}=\{n\in\mathbb{N}:|a_{n}-1|<\epsilon\}\in\mathscr{U}$ and for all $j>1$, $X_{j}=\{n\geq j:|a_{n-j}|<\epsilon\}\in\mathscr{U}$. However, $X_{j}=(X_{1}+j)\setminus j$, so by idempotency, $\{j\in\mathbb{N}:X_{1}+j\in\mathscr{U}\}\in\mathscr{U}$, so $\varphi(\pi(s^{j})^{*}a\pi(s^{j}))=1$ for some $j\in\mathbb{N}$. Therefore, there is no $a\in\mathscr{Q}(H)$ s.t. $\pi_{\varphi}(a)\xi_{j}=\delta_{j0}\xi_{j}$, so infinite transitivity fails. ∎

### v4

@@ 4. Existence of P-states
Let $(q_{j})$ be a scalar sub-$\sigma$ sequence. As $q_{j}$ is a scalar projection, there is a unique $\varphi_{j}\in P(A)$ such that
$\mathscr{F}_{\varphi_{j}}=\{a\in A_{+,1}:q_{j}\leq a\}.$
We define $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ as in 4.3. By 4.3, if $\mathscr{U}$ is not a P-point, $\varphi_{\mathscr{U}}$ is not a P-state
Conversely, suppose $\mathscr{U}$ is a P-point. By the scalar sub-$\sigma$ property, fix a faithful, nondegenerate representation $\rho:A\rightarrow\mathscr{B}(H)$ such that $\rho(q_{j})$ is finite dimensional. There is a unique extension $\tilde{\rho}:\mathscr{M}(A)\rightarrow\mathscr{B}(H)$, per [FA19, 13.2.1]. Fix an orthonormal basis $(\xi_{\gamma}:\gamma<\kappa)$ of $H$ which simultaneously diagonalizes $\tilde{\rho}(q_{j})$ for every $j$ and express $\tilde{\rho}(q_{j})=\text{proj}_{\overline{\text{span}}\{\xi_{\gamma}:\gamma\in\Lambda_{j}\}}$ for some $\Lambda_{j}\Subset\kappa$.
Fix $a\in\mathscr{F}_{\varphi_{\mathscr{U}}}$ and a lift $\tilde{a}\in\mathscr{M}(A)$ such that $\|\tilde{a}\|=1$. Now let
$X_{n}=\big\{k\in\mathbb{N}:\tilde{\varphi_{k}}(\tilde{a})>1-\frac{1}{n}\big\}.$
Since $\mathscr{U}$ is a P-point, there is $X\in\mathscr{U}$ such that $X\subseteq^{*}X_{n}$. We first claim that
$\|\pi(\rho(p_{X}(1-\tilde{a})p_{X}-\sum_{j\in{X}}q_{j}(1-\tilde{a})q_{j}))\|=0,$
where $\pi:\mathscr{B}(H)\rightarrow\mathscr{Q}(H)$ is the quotient map. Suppose this is not the case. Then there exists $\epsilon>0$ and infinitely many $\gamma<\kappa$ such that
$\|\tilde{\rho}\big(\sum_{\begin{subarray}{c}i,j\in X\\ i\neq j\end{subarray}}q_{i}\tilde{a}q_{j}\big)\xi_{\gamma}\|\geq\epsilon.$
Some linear algebra yields
$\|\tilde{\rho}\big(\sum_{\begin{subarray}{c}i,j\in X\\ i\neq j\end{subarray}}q_{i}\tilde{a}q_{j}\big)\xi_{\gamma}\|^{2}=\sum_{\begin{subarray}{c}i\in X\\ i\neq j\end{subarray}}\|\tilde{\rho}(q_{i}\tilde{a}q_{j})\xi_{\gamma}\|^{2},$
when $\gamma\in\Lambda_{j}$ and 0 otherwise. Now,
$\|\tilde{\rho}(p_{X}\tilde{a}p_{X})\xi_{\gamma}\|^{2}$ $=\sum_{i\in X}\|\tilde{\rho}(q_{i}\tilde{a}q_{j})\xi_{\gamma}\|^{2}$
$\geq\|\tilde{\rho}(q_{j}\tilde{a}q_{j})\xi_{\gamma}\|^{2}+\epsilon^{2}$
$\geq\tilde{\varphi_{j}}(\tilde{a})^{2}\|\tilde{\rho}(q_{j})\xi_{\gamma}\|^{2}+\epsilon^{2}$
$\geq\tilde{\varphi_{j}}(\tilde{a})^{2}+\epsilon^{2}$
when $\gamma\in\Lambda_{j}$ as in the assumption. Since there are infinitely many such $\gamma$, we can take $\gamma\in\Lambda_{j}$ with $j\in X\setminus k$ for arbitrarily large $k$,
$\|\tilde{a}\|^{2}\geq\|\tilde{\rho}(p_{X}\tilde{a}p_{X})\|^{2}\geq\|\tilde{\rho}(p_{X}\tilde{a}p_{X})\xi_{\gamma}\|^{2}\geq\big(1-\frac{1}{n}\big)^{2}+\epsilon^{2}>1$
for sufficiently large $n$, a contradiction. Thus, $\tilde{\rho}(p_{X}(1-\tilde{a})p_{X}-\sum_{j\in X}q_{j}(1-\tilde{a})q_{j})\in\mathscr{K}(H)$. Now, because $e_{k}\rightarrow 1_{\mathscr{B}(H)}$ strongly, $p_{X}(1-\tilde{a})p_{X}-\sum_{j\in X}q_{j}(1-\tilde{a})q_{j}\in A$, so back in $\mathscr{M}(A)$ with $\pi:\mathscr{M}(A)\rightarrow\mathscr{Q}(A)$ the quotient map
$\|\pi(p_{X}(1-\tilde{a})p_{X})\|$ $=\|\pi(\sum_{j\in X}q_{j}(1-\tilde{a})q_{j})\|$
$=\lim_{k\rightarrow\infty}\|(1-e_{k})\sum_{j\in X}q_{j}(1-\tilde{a})q_{j}\|$
$=\lim_{k\rightarrow\infty}\|\sum_{j\in X\setminus k}q_{j}(1-\tilde{a})q_{j}\|$
$=\lim_{k\rightarrow\infty}\sup_{j\in X\setminus k}\|q_{j}(1-\tilde{a})q_{j}\|$
$=\lim_{k\rightarrow\infty}\sup_{j\in X\setminus k}\tilde{\varphi_{j}}(1-\tilde{a})=0.$
Therefore,
$\|(1-a)\pi(p_{X})\|^{2}$ $\leq\|(1-a)^{\frac{1}{2}}\pi(p_{x})\|^{2}$
$\leq\|\pi(p_{X}(1-\tilde{a})p_{X})\|=0,$
so $\pi(p_{X})\ll a$.
Finally, fix $(a_{n})$ in $\mathscr{F}_{\varphi_{\mathscr{U}}}$ and $X_{n}\in\mathscr{U}$ so that $\pi(p_{X_{n}})\ll a_{n}$. Since $\mathscr{U}$ is a P-point, there is $X\in\mathscr{U}$ with $X\subseteq^{*}X_{n}$. Therefore $\pi(p_{X})\ll\pi(p_{X_{n}})\ll a_{n}$. ∎
@@
Let $A$ be nonunital, $\sigma$-unital with a sub-$\sigma$ sequence, and $\varphi\in P(\mathscr{Q}(A))$. If $\varphi$ is a P-state, then $\mathscr{X}_{\varphi}$ is a non-meagre P-filter. Also, for any decreasing $X_{n}\subseteq\mathbb{N}$, there exists a pseudointersection $X$ of $(X_{n})$ such that $\varphi(\pi(p_{X}))=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Also, if $\varphi(\pi(p_{X_{n}}))\not\rightarrow 0$, there exists $(k_{n})$ in $\mathbb{N}$ increasing and $X\subseteq\mathbb{N}$ such that $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ and $\varphi(X)>0$.
@@
Using the obvious type, there exists $a\in\mathscr{P}_{\varphi}$ such that $a\ll\pi(p_{X_{n}})$ and $\varphi(a)=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Then by 4.2, there is a pseudointersection $X$ of $(X_{n})$ such that $a\ll\pi(p_{X})$, so $1=\varphi(a)\leq\varphi(\pi(p_{X}))$. It then follows that $\mathscr{X}_{\varphi}$ is a P-filter.
@@
for $i<l$, where $k<m_{0}<\cdots<m_{l}$, then $l\epsilon<3^{n+1}$. We may then choose $k_{n}$ such that $\varphi(\pi(p_{X_{j}}))>\epsilon$ for some $k_{n}\leq j<k_{n+1}$ and
@@
By the normality ensured in the second paragraph, we have $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ so that $\varphi(\pi(p_{X}))\geq\frac{\epsilon}{2}>0$. ∎
@@
    For every nonunital, $\sigma$-unital $A$ admitting a scalar sub-$\sigma$ sequence, there exist $2^{\mathfrak{c}}$-many inequivalent P-states $\varphi$ on $\mathscr{Q}(A)$.
@@
    For every nonunital, $\sigma$-unital $A$ admitting a sub-$\sigma$ sequence, there exist no P-states on $\mathscr{Q}(A)$.
@@
For (1), under the continuum hypothesis, there are $2^{\mathfrak{c}}$ RK-inequivalent selective ultrafilters (which are P-points) on $\mathbb{N}$, so the map $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ in 4.4 produces $2^{\mathfrak{c}}$ inequivalent P-states.
For (2), we use the forcing argument outlined in lemma [BI13, 6.5], where we use 4.5 in place of [BI13, 6.2]. In this model, there are no $\varphi$ on $\mathscr{Q}(A)$ having the conclusion of 4.5, so there are no P-states. ∎
@@ 4.1. Counterexample to Infinite Transitivity
The question of whether infinite transitivity in $(\mathscr{Q}(H),\pi_{\varphi})$ implies degree-1 saturation of $(\mathscr{Q}(H),\varphi)$ is fairly opaque. In this section, we at least prove that in ZFC, the Calkin algebra has an irreducible representation which does not admit infinite transitivity.
@@
For $X\subseteq\mathbb{N}$, $j\in\mathbb{N}$, we define $X+j=\{m+j:m\in X\}$ and for an ultrafilter $\mathscr{U}$ on $\mathbb{N}$, $\mathscr{U}+j=\{X+j:X\in\mathscr{U}\}$. Since $\beta\mathbb{N}$ is compact, we can define $\mathscr{U}+\mathscr{V}=\lim_{j\rightarrow\mathscr{V}}(\mathscr{U}+j)$. We call a nonprincipal ultrafilter $\mathscr{U}$ idempotent if $\mathscr{U}+\mathscr{U}=\mathscr{U}$.
By [EL58, 1], every compact topological space with a semigroup structure whose left operation is continuous admits an idempotent element. The addition defined in 4.7 plainly grants this semigroup structure on $\beta\mathbb{N}$, so there is $\mathscr{U}\in\beta\mathbb{N}$ idempotent, and this cannot be principal.
@@
Let $\mathscr{U}$ be an idempotent ultrafilter on $\mathbb{N}$. We claim that $\pi=\pi_{\varphi_{\mathscr{U}}}$ fails infinite transitivity. Let $s\in\mathscr{B}(H)$ be the left shift along an orthonormal basis $(\eta_{n})$ of $H$ and $\varphi_{\mathscr{U}}$ as in 4.4 along $(\eta_{n})$. Let $\xi_{j}=\pi(s^{j})+L_{\varphi_{\mathscr{U}}}$. Then the equations
$\pi(a)\xi_{j}=\delta_{j0}\xi_{j}$
@@
where $a_{n}=\langle a\eta_{n},\eta_{n}\rangle$. Thus $X_{1}=\{n\in\mathbb{N}:|a_{n}-1|<\epsilon\}\in\mathscr{U}$ and for all $j>1$, $X_{j}=\{n\geq j:|a_{n-j}|<\epsilon\}\in\mathscr{U}$. However, $X_{j}=(X_{1}+j)\setminus j$, so by idempotency, $\{j\in\mathbb{N}:X_{1}+j\in\mathscr{U}\}\in\mathscr{U}$, so $\varphi(\pi(s^{j})^{*}a\pi(s^{j}))=1$ for some $j\in\mathbb{N}$. Therefore, there is no $a\in\mathscr{Q}(H)$ such that $\pi(a)\xi_{j}=\delta_{j0}\xi_{j}$, so infinite transitivity fails. ∎